Find the conditional probability of the indicated event when two fair dice (one red and one green) are rolled. HINT [See Example 1.]

The sum is 4, given that the green one is either 1 or 2.
Answer #1

Sum 4 can be obtained in 3 ways i.e. (3, 1), (1, 3), (2, 2)

In this, there would be two ways when green one will be at 1 or 2. Also,

Total number of outcomes with green one at 1 or 2 = 6 + 6 = 12

Hence,

P(Sum is 4 | Green one is either 1 or 2) = 2/12 = 1/6 = 0.1667